### v1

@@ 1 Introduction
One such mechanism is cosmological gravitational particle production (GPP) Parker:1969au ; Parker:1971pt ; Ford:2021syk ; Chung:1998zb ; Chung:1998ua in which particles are produced during inflation due to the rapid expansion of the universe. Particles produced via GPP are generally thought to be in the supermassive range (i.e., $m\gtrsim H_{e}$ where $H_{e}$ is the Hubble parameter at the end of inflation), but can also have masses in the ultralight regime Kolb:2020fwh ; Graham:2015rva ; Ahmed:2020fhc . Previous work has shown that GPP is a viable production mechanism for particles with a wide variety of masses and spins Chung:1998zb ; Graham:2015rva ; Kolb:2020fwh ; Alexander:2020gmv ; Kolb:2021xfn ; Kolb:2021nob ; Kolb:2023dzp ; Capanelli:2023uwv ; Kaneta:2023uwi ; Maleknejad:2022gyf and in both single field and multifield inflation models Kolb:2022eyn . In particular, Refs. Graham:2015rva ; Kolb:2020fwh have shown that spin-1, or ‘dark photon’ dark matter can be gravitationally produced to yield the correct DM relic density. Dark photons have been considered as a portal from the standard model to a dark sector, but can also be considered a DM candidate in their own right, see e.g., Redondo_2009 ; Nelson:2011sf ; Long:2019lwl ; Arias:2012az ; Graham:2015rva ; Kolb:2020fwh ; Co:2018lka ; Dror:2018pdh ; Bastero-Gil:2018uel ; Bastero-Gil:2021wsf ; Agrawal:2018vin ; Krnjaic:2022wor ; Pospelov:2008jk ; Fradette_2014 ; Jaeckel_2013 ; East:2022rsi ; Cyncynates:2023zwj .
Previous work on GPP of dark photon dark matter has largely considered particles which are minimally coupled to gravity Graham:2015rva ; Kolb:2020fwh . However, a theory of a massive spin-1 field can also contain nonminimal couplings to gravity; general principles of effectice field theory (EFT) demand their inclusion. These additional couplings are an important factor in fully understanding gravitational production of spin-1 particles and dark photon dark matter. Recent works have begun to discuss these possibilities Capanelli:2023uwv ; Ozsoy:2023gnl ; Cembranos:2023qph , however a full survey of the viable parameter space while also taking into account instabilities in the theory is yet to appear in the literature.

### v2

@@ 1 Introduction
One such mechanism is cosmological gravitational particle production (GPP) Parker:1969au ; Parker:1971pt ; Ford:2021syk ; Chung:1998zb ; Chung:1998ua in which particles are produced during inflation due to the rapid expansion of the universe. Particles produced via GPP are generally thought to be in the supermassive range (i.e., $m\gtrsim H_{e}$ where $H_{e}$ is the Hubble parameter at the end of inflation), but can also have masses in the ultralight regime Kolb:2020fwh ; Graham:2015rva ; Ahmed:2020fhc . Previous work has shown that GPP is a viable production mechanism for particles with a wide variety of masses and spins Chung:1998zb ; Graham:2015rva ; Ema:2015dka ; Ema:2016hlw ; Ema:2019yrd ; Kolb:2020fwh ; Ahmed:2020fhc ; Alexander:2020gmv ; Kolb:2021xfn ; Kolb:2021nob ; Kolb:2023dzp ; Capanelli:2023uwv ; Kaneta:2023uwi ; Maleknejad:2022gyf and in both single field and multifield inflation models Kolb:2022eyn . In particular, Refs. Graham:2015rva ; Kolb:2020fwh have shown that spin-1, or ‘dark photon’ dark matter can be gravitationally produced to yield the correct DM relic density. Dark photons have been considered as a portal from the standard model to a dark sector, but can also be considered a DM candidate in their own right, see e.g., Redondo_2009 ; Nelson:2011sf ; Long:2019lwl ; Arias:2012az ; Graham:2015rva ; Kolb:2020fwh ; Co:2018lka ; Dror:2018pdh ; Bastero-Gil:2018uel ; Bastero-Gil:2021wsf ; Agrawal:2018vin ; Krnjaic:2022wor ; Pospelov:2008jk ; Fradette_2014 ; Jaeckel_2013 ; Nakayama:2019rhg ; East:2022rsi ; Cyncynates:2023zwj ; Zhang:2023fhs ; Kitajima:2023fun ; Bhattacharyya:2023kbh .
Previous work on GPP of dark photon dark matter has largely considered particles which are minimally coupled to gravity Graham:2015rva ; Ema:2019yrd ; Kolb:2020fwh ; Ahmed:2020fhc . However, a theory of a massive spin-1 field can also contain nonminimal couplings to gravity; general principles of effectice field theory (EFT) demand their inclusion. These additional couplings are an important factor in fully understanding gravitational production of spin-1 particles and dark photon dark matter. Recent works have begun to discuss these possibilities Alonso-Alvarez:2019ixv ; Capanelli:2023uwv ; Ozsoy:2023gnl ; Cembranos:2023qph , however a full survey of the viable parameter space while also taking into account instabilities in the theory is yet to appear in the literature.
@@ 4.3 Superluminal Propagation
There are of course theories in which superluminal propagation of perturbations does not lead to causality paradoxes, such as k-essence models Babichev:2007dw or noncommutative geometries Hashimoto:2000ys . However, there are others in which superluminal propagation is indeed causally problematic (e.g. Adams:2006sv ; DeFelice:2006pg ). A full analysis of the causality structure of the theory is beyond the scope of our current analysis, and so to be conservative, we choose to impose (sub)luminality restrictions on the sound speed. Even with this restriction, the theory admits a broad parameter space for a viable cosmological relic density of dark photons.